A boy tosses a ball straight up with an initial velocity of 13.0m/s a)how long does the ball take to reach its highest point? B)how high does the ball rise above it's realise point


1
Expert's answer
2021-03-09T07:46:53-0500

Given,

Initial velocity of the ball (u)=13 m/s

Let the time taken by the ball to reach to highest point be t and the height reached by the ball be H.

At the top most point, the velocity of the ball will be zero, so "v=0"

Gravitational acceleration of the ball "(g)=9.8 m\/s^2"

Now, applying the first law of motion,

"v=u-gt"

Now, substituting the values,

"0=13-9.8t"

"\Rightarrow t=\frac{13}{9.8}sec"

Now, applying the third equation of motion,

"v^2=u^2-2gh"

Substituting the values,

"0=13^2-2\times 9.8\times h"


"\Rightarrow h = \frac{13\times 13}{2\times 9.8}m"

"h=8.62m"